OpenStudy (anonymous):

cot(cos^-1u) please help I just don't understand these questions!!

ganeshie8 (ganeshie8):

use the identity : \(\large 1 + \tan^2x = \sec^2x\) put \(x = \cos^{-1}u\)

ganeshie8 (ganeshie8):

\(\large 1 + \tan^2(\cos^{-1}u) = \frac{1}{\cos^2(\cos^{-1}u)}\) \(\large 1 + \tan^2(\cos^{-1}u) = \frac{1}{u^2}\) \(\large \tan^2(\cos^{-1}u) = \frac{1-u^2}{u^2}\) \(\large \cot (\cos^{-1}u) = \frac{u}{\sqrt{1-u^2}}\)

OpenStudy (anonymous):

Thank you so much for your help!! Trig is one confusing subject. Can you tell me what you did with tan^2 in the second to last part? No really understanding that part.

ganeshie8 (ganeshie8):

very good question :) u knw tan = 1/cot right ?

ganeshie8 (ganeshie8):

\(\large \tan^2(\cos^{-1}u) = \frac{1-u^2}{u^2}\) \(\large \frac{1}{\tan^2(\cos^{-1}u)} = \frac{u^2}{1-u^2}\) \(\large \cot^2(\cos^{-1}u) = \frac{u^2}{1-u^2}\)

ganeshie8 (ganeshie8):

take square-root both sides

OpenStudy (anonymous):

oh thank you for explaining. It makes so much more sense now. I should never take a online math class again.

ganeshie8 (ganeshie8):

glad it made some sense :)